# Mitigating Sensitive Information Leakage in LLMs4Code through Machine Unlearning

### v1

@@ Abstract
Large Language Models for Code (LLMs4Code) excel at code generation tasks, yielding promise to release developers from huge software development burdens. Nonetheless, these models have been shown to suffer from the significant privacy risks due to the potential leakage of sensitive information embedded during training, known as the memorization problem. Addressing this issue is crucial for ensuring privacy compliance and upholding user trust, but till now there is a dearth of dedicated studies in the literature that focus on this specific direction. Recently, machine unlearning has emerged as a promising solution by enabling models to “forget” sensitive information without full retraining, offering an efficient and scalable approach compared to traditional data cleaning methods. In this paper, we empirically evaluate the effectiveness of unlearning techniques for addressing privacy concerns in LLMs4Code. Specifically, we investigate three state-of-the-art unlearning algorithms and three well-known open-sourced LLMs4Code, on a benchmark that takes into consideration both the privacy data to be forgotten as well as the code generation capabilites of these models. Results show that it is feasible to mitigate the privacy concerns of LLMs4Code through machine unlearning while maintain their code generation capabilities at the same time. We also dissect the forms of privacy protection/leakage after unlearning and observe that there is a shift from direct leakage to indirect leakage, which underscores the need for future studies addressing this risk.
@@ 1 Introduction
In recent, the remarkable success of Large language models (LLMs) (Brown, 2020) in diverse natural language processing tasks has been tailored for the realm of software engineering, yielding several language models tailored specifically for code, termed as Large Language Models for Code (LLMs4Code), e.g., CodeLlama and Stable Code (Xu et al., 2022). With extensive pre-training process on programming language datasets, such models have demonstrated proficiency with prominent performance on code-related tasks (Geng et al., 2024; Deng et al., 2024; Qin et al., 2024). For instance, Geng et al. (2024) has demonstrated the capability of LLMs4Code in producing code summarizations that are not only of superior quality but also cater to the varied requirements of human programmers through in-context learning(Geng et al., 2024).
As a double-edged sword of LLMs4Code, the risk of sensitive information including Personally Identifiable Information (PII), private data, or confidential secrets has already been highlighted by recent studies (Yang et al., 2024b; Jahanshahi and Mockus, 2025). From the perspective of model attacking (Yao et al., 2024; Dong et al., 2024), empirical evidence reveal that specific prompts can result in the leakage of corresponding sensitive information (Huang et al., 2024; Carlini et al., 2021). More formally, such risk of privacy disclosures during the utilization of LLMs4Code is commonly termed as the memorization problem (Al-Kaswan and Izadi, 2023; Lukas et al., 2023). Since the memorization problem commonly exists in a wide range of code-relevant tasks, e.g., code generation (Svyatkovskiy et al., 2020; Wang et al., 2023a), and yield inevitable risk of developers in their daily development activities, we argue that:
@@
However, to the best of our knowledge, concurrent work rarely provides solutions to such an important but challenging problem, offering only empirical findings on the memorization problem (Leybzon and Kervadec, 2024; Kiyomaru et al., 2024). We also note that one potential approach involves incorporating a dedicated data cleaning phase when pre-processing the training data, yielding the inflexibility and unscalability due to the necessitate of substantial engineering endeavors to devise appropriate rules and heuristics.
Fortunately, Machine Unlearning (MU) has emerged as a technique striving to assist the target model in “forgetting” the data points from the initial training set, offering lightweight but effective approach to aid in protecting sensitive information from LLMs (Liu et al., 2024; Nguyen et al., 2022). To be specific, MU proposes to resemble an “auxiliary” sanitized dataset (devoid of sensitive information), thereby potentially saving considerable development costs compared to retraining the model from scratch. Following such an intuition, a number of studies have verified the promising effectiveness of MU on making LLMs forget specific contents that they met during training (Chen and Yang, 2023; Chundawat et al., 2023). Nonetheless, we also note that the current literature lacks a comprehensive understanding regarding the strengths and weaknesses of existing MU techniques within the context of LLMs4Code, including their effectiveness on mitigating the privacy leakage during the code generation process, and how the state-of-the-art MU techniques will perform on LLMs4Code.
@@
    MU is a promising way to simultaneously mitigate the privacy concerns of LLMs4Code while maintain their code generation capabilities at the same time. Specifically, unlearning can decrease the leak rate of AIXCoder by more than 50% while only bring a negalegable side effect to code generation.
@@
    After unlearning, LLMs4Code become more likely to leak the privacy in an indirect manner, which means they tend to leak the information that is not explicitly queried. This suggests that future works should also take into consideration the indirect privacy leakage for a more robust unlearning process.
@@ 2.1 LLMs & LLMs4Code
LLMs and LLMs4Code are significant innovations in the fields of natural language processing and programming. LLMs, exemplified by models like ChatGPT, are trained on massive text datasets to comprehend and generate human language with remarkable accuracy and fluency. These models have demonstrated capabilities in a wide array of language-related tasks, from translation and summarization to dialogue generation and content creation (Ugare et al., 2024; Feng et al., 2024; Yang et al., 2024a). On the other hand, LLMs4Code, such as OpenAI’s Codex and GitHub’s Copilot, are specialized variants tailored for programming tasks. By integrating knowledge of both natural language and programming languages, LLMs4Code excel in assisting developers by providing code suggestions (Dong et al., 2023; Ahmed et al., 2024), auto-completions(Li et al., 2024), and even entire code segments (Zhang et al., 2023; Wang et al., 2023b; Dong et al., ), elevating efficiency and creativity in software development processes.
@@ 2.2 Memorization Problem
Numerous studies have presented the evidence of the susceptibility of LLMs to retaining training data. A typical way to observe such a phenomenon is to perform extraction attacks. For instance, Carlini et al. (Carlini et al., 2021) illustratively extracted hundreds of verbatim text sequences, including sensitive personal identifiers, from GPT-2’s training corpus. The modus operandi of their attack entailed crafting a series of prompts and subsequently assessing whether the sequences generated in response to these prompts were present within the training dataset. In another study, Liang Niu’s work (Niu et al., 2023) further validated the efficacy of such an assault strategy, introducing the concept of perplexity as a metric to gauge the model’s degree of surprise regarding the generated sequences. In this context, a lower perplexity value indicates a higher likelihood that the sequence has been encountered during training, thereby suggesting familiarity on the part of the model. Formally, perplexity is quantified as follows:
@@
where $log\quad f(x_{i}|x_{1},\ldots,x_{i-1},\Theta)$ indicates the log likelihood of the token $x_{i}$ given all previous tokens $x_{1},...,x_{i-1}$. Besides, a straightforward heuristic based on knowledge distillation can completely expunge the information that needs to be forgotten, while preserving an accuracy exceeding 90% on the retained data (Chundawat et al., 2023).
The aforementioned studies collectively highlight that addressing the memorization problem of LLMs is crucial for safeguarding user privacy. Very recently, a latest study empirically demonstrates that memorization problem also occurs in LLMs4Code, showcasing that hard-coded credentials in code can be easily leaked during the code completion process (Huang et al., 2024). Therefore, our study aims to understand how well the problem can be mitigated, which builds the foundation for mitigating the potential risks associated with the retention of sensitive information stored in LLMs4Code.
@@ 2.3 Machine Unlearning
Relevant to the issue of privacy leaks in LLMs (Huang et al., 2022), the concept of unlearning involves recalibrating a model after training to erase certain contents from its captured knowledge (Jagielski et al., 2020; Jang et al., 2022), thereby avoiding the costly process of retraining LLMs. Given the ability of large language models to recall specific details from their training sets, the targeted deletion of such privacy-sensitive data is of significant value.
Pioneer efforts by Chen & Yang (2023) (Chen and Yang, 2023) and Eldan & Russinovich (2023) (Eldan and Russinovich, 2023) have provided model designers with post-training modifications that can protect privacy at relatively low computational costs. Nevertheless, unlearning algorithms are still at an early stage of development, with different methods showing varying degrees of effectiveness and no definitive benchmark for evaluating their performance. To adequately assess the effectiveness of unlearning, a comprehensive evaluation framework is essential. This framework should include metrics that not only quantify the reduction of information related to the target data, but also evaluate the functional behaviour of the model after unlearning, ensuring that the overall performance of the model remains intact, while effectively omitting sensitive information. Our study aims at building such a benchmark for machine unlearning on LLMs4Code. Through rigorous evaluations on such a benchmark, we can move towards a more comprehensive understanding of the capabilities of machine unlearning and its significance in enhancing privacy-preserving practices within the application of LLMs4Code.
Figure 1: The Workflow of Our Study.
3 Experiment Settings
Figure 1 demonstrates the workflow of this study. Suppose that LLMs4Code can initially work well on code generation (given the general query) but leak privacy information when given a specific code completion prompt (i.e., the privacy query). Our work aims at investigating that after applying several machine unlearning techniques to the models, whether they would (1) avoid sensitive information leakage give nthe privacy query, and (2) preserve the capability for code generation given the general query.
@@ 3.1 Research Questions
  • RQ3: Privacy Leakage Forms after Unlearning This question investigates how the sensitive information would be still leaked after unlearning. Especially, we focus on analyzing if the leakage is intentional or unintentional.
@@ 3.2 Dataset
As for the forget set, we followed the previous study (Maini et al., 2024) and created a synthetic dataset consisting of 500 fictional character resumes. Each resume contains essential details such as name, address, education, phone number, and email address. The sensitive attributes in this study is categorized as follows:
@@
To evaluate the potential privacy risks posed by LLMs4Code, we designed 10 privacy-related questions for each fictional character. These questions are tailored to probe various aspects of personal data security and privacy concerns, including sensitive attributes like bank accounts and other distinct properties. In order to alighn with the main functionality of LLMs4Code (i.e., code generation), these questions are transformed into the format of code completions (a detailed example is listed in the Appendix). Furthermore, these 5K questions are randomly split into train and test sets with a ration of 9:1, where the train set is used to perform the unlearning while the test set is used to evlaute the performance of the unlearning.
As for the retain set, we leveraged the well-known established datasets HumanEval (Chen et al., 2021), MBXP (Athiwaratkun et al., 2022) and LiveBenchWhite et al. (2024) in the code generation domain. Similar to the forget set, we randomly collected 5K samples from these two datasets and split them into train and test sets following a standard 9:1 ratio.
@@ 3.3 Unlearning Techniques
Forget Set Gradient Ascent (GA) (Liu et al., 2022): The key idea of this approach is to calculate the gradients of the loss function with respect to the model’s parameters based on the responses to the privacy queries in the forget set. The gradients are then used to update the model’s parameters in a direction that encourages the model to “forget” the patterns associated with answering privacy questions directly. Formally,
@@
Forget Set Gradient Ascent and Retain Set Gradient Descent (GA+GD) (Liu et al., 2022): The key idea of this approach is to simultaneously perform the forget set gradient ascent along with the retain set gradient descent, aiming at retaining the models’ capabilities to handle general queries. Formally,
@@
where $S_{R}$ denotes the retain set, $L(\S_{R},w)$ is the loss function for the retain set $S_{R}$, and $L_{\text{diff}}$ represents the combined effect on the model’s loss.
Forget Set Gradient Ascent and Kullback-Leibler Divergence Minimization (GA+KL) (Rafailov et al., 2024): This approach provides another way to simutaneously forget the privacy-sensitive information and maintain its proficiency in handling general queries. The key idea is to minimize the difference between the current model’s distribution of responses and the desired distribution after the unlearning process. To that end, the KL divergence between the model’s responses and the responses of a model that has been fine-tuned on the retain set is calculated. Formally,
@@
where $\mathrm{KL}$ measures the difference between the output distributions of the fine-tuned model $M_{\mathrm{fin}}$ and the unlearning model $M_{\mathrm{unl}}$ for subsequences $s_{<i}$, the summations over $s\in S_{R}$ and $i$ are used to calculate the KL divergence, and $L_{\mathrm{KL}}$ represents the overall loss for the KL divergence-based unlearning process.

### v2

@@ Abstract
Large Language Models for Code (LLMs4Code) have achieved strong performance in code generation, but recent studies reveal that they may memorize and leak sensitive information contained in training data, posing serious privacy risks. To address this gap, this work presents the first comprehensive empirical study on applying machine unlearning to mitigate sensitive information leakage in LLMs4Code. We first construct a dedicated benchmark that includes: (i) a synthetic forget set containing diverse forms of personal information, and (ii) a retain set designed to evaluate whether code-generation capability is preserved after unlearning. Using this benchmark, we systematically assess three representative unlearning algorithms (GA, GA+GD, GA+KL) across three widely used open-source LLMs4Code models (AIXCoder-7B, CodeLlama-7B, CodeQwen-7B). Experimental results demonstrate that machine unlearning can substantially reduce direct memorization-based leakage: on average, the direct leak rate drops by more than 50% while retaining about over 91% of the original code-generation performance. Moreover, by analyzing post-unlearning outputs, we uncover a consistent shift from direct to indirect leakage, revealing an underexplored vulnerability that persists even when the target data has been successfully forgotten. Our findings show that machine unlearning is a feasible and effective solution for enhancing privacy protection in LLMs4Code, while also highlighting the need for future techniques capable of mitigating both direct and indirect leakage simultaneously.
keywords:
[…]
{highlights}
Machine unlearning is a promising way to simultaneously mitigate the privacy concerns of LLMs4Code while maintaining their code generation capabilities at the same time. Specifically, unlearning can decrease the leak rate of AIXCoder by more than 50% while only bringing a negligible side effect to code generation.
After unlearning, LLMs4Code become more likely to leak the privacy indirectly, which means they tend to leak the information that is not explicitly queried. This suggests that future works should also take into consideration the indirect privacy leakage for a more robust unlearning process.
All code and data in this study are publicly available at https://doi.org/10.5281/zenodo.14729266.
@@ 1 Introduction
Recently, the remarkable success of Large language models (LLMs) (brown2020language; qiu2025dcsolver; xu2025efficient) in diverse natural language processing tasks has been tailored for the realm of software engineering, yielding several language models tailored specifically for code, termed as Large Language Models for Code (LLMs4Code), e.g., CodeLlama and Stable Code (xu2022systematic). With an extensive pre-training process on programming language datasets, such models have demonstrated proficiency with prominent performance on code-related tasks (geng2024large; deng2024large; qin2024agentfl). For instance, geng2024large has demonstrated the capability of LLMs4Code in producing code summarizations that are not only of superior quality but also cater to the varied requirements of human programmers through in-context learning(geng2024large).
As a double-edged sword of LLMs4Code, the risk of sensitive information, including Personally Identifiable Information (PII), private data, or confidential secrets, has already been highlighted by recent studies (yang2024robustness; jahanshahi2025cracks). From the perspective of model attacking (yao2024survey; dong2024attacks), empirical evidence reveals that specific prompts can result in the leakage of corresponding sensitive information (huang2024your; carlini2021extracting). More formally, such risk of privacy disclosures during the utilization of LLMs4Code is commonly termed as the memorization problem (al2023ab; lukas2023analyzing). Since the memorization problem commonly exists in a wide range of code-relevant tasks, e.g., code generation (svyatkovskiy2020intellicode; wang2023codet5+), and yields an inevitable risk to developers in their daily development activities, we argue that:
@@
However, to the best of our knowledge, concurrent work rarely provides solutions to such an important but challenging problem, offering only empirical findings on the memorization problem (leybzon2024learning; kiyomaru2024comprehensive). We also note that one potential approach involves incorporating a dedicated data cleaning phase during pre-processing of the training data, which yields inflexibility and unscalability due to the necessity of substantial engineering efforts to devise appropriate rules and heuristics.
Fortunately, Machine Unlearning (MU) has emerged as a technique striving to assist the target model in “forgetting” the data points from the initial training set, offering a lightweight but effective approach to aid in protecting sensitive information from LLMs (liu2024machine; nguyen2022survey). To be specific, MU proposes to resemble an “auxiliary” sanitized dataset (devoid of sensitive information), thereby potentially saving considerable development costs compared to retraining the model from scratch. Following such an intuition, a number of studies have verified the promising effectiveness of MU on making LLMs forget specific contents that they met during training (chen2023unlearn; chundawat2023zero). Nonetheless, we also note that the current literature lacks a comprehensive understanding regarding the strengths and weaknesses of existing MU techniques within the context of LLMs4Code, including their effectiveness on mitigating the privacy leakage during the code generation process, and how the state-of-the-art MU techniques will perform on LLMs4Code. While several prior works (e.g., chen2023unlearn; chundawat2023zero) have explored machine unlearning techniques on general-purpose LLMs, applying such methods directly to LLMs4Code is non-trivial. Code generation models differ significantly from natural language models in that code is more repetitive, memorization-prone, and functionally precise. These properties make unlearning in LLMs4Code both technically more challenging and practically more consequential. Furthermore, the evaluation of unlearning effectiveness in general LLMs often relies on classification accuracy or perplexity, which do not apply to code generation tasks where pass@1 and functional correctness are the primary metrics. Our work addresses this gap by systematically evaluating unlearning strategies on LLMs4Code using both privacy leakage and code generation metrics.
@@
    MU is a promising way to simultaneously mitigate the privacy concerns of LLMs4Code while maintaining their code generation capabilities at the same time. Specifically, unlearning can decrease the leak rate of AIXCoder by more than 50% while only bringing a negligible side effect to code generation.
@@
    After unlearning, LLMs4Code become more likely to leak the privacy indirectly, which means they tend to leak the information that is not explicitly queried. This suggests that future works should also take into consideration the indirect privacy leakage for a more robust unlearning process.
  4. 4.
    All code and data in this study are publicly available at https://doi.org/10.5281/zenodo.14729266.
@@ 2.1 LLMs & LLMs4Code
LLMs and LLMs4Code are significant innovations in the fields of natural language processing and programming. LLMs, exemplified by models like ChatGPT, are trained on massive text datasets to comprehend and generate human language with remarkable accuracy and fluency. These models have demonstrated capabilities in a wide array of language-related tasks, from translation and summarization to dialogue generation and content creation (ugare2024improving; feng2024improving; yang2024exploring). On the other hand, LLMs4Code, such as OpenAI’s Codex and GitHub’s Copilot, are specialized variants tailored for programming tasks. By integrating knowledge of both natural language and programming languages, LLMs4Code excel in assisting developers by providing code suggestions (dong2023codescore; ahmed2024automatic), auto-completions(li2024ircoco), and even entire code segments (zhang2023toolcoder; wang2023chatcoder; dong2304self), elevating efficiency and creativity in software development processes.
@@ 2.2 Memorization Problem
Numerous studies have presented evidence of the susceptibility of LLMs to retaining training data. A typical way to observe such a phenomenon is to perform extraction attacks. For instance, Carlini et al. (carlini2021extracting) illustratively extracted hundreds of verbatim text sequences, including sensitive personal identifiers, from GPT-2’s training corpus. Their attack strategy involved crafting a series of prompts and then evaluating whether the sequences generated in response appeared in the training dataset. In another study, Liang Niu’s work (niu2023codexleaks) further validated the efficacy of such an assault strategy, introducing the concept of perplexity as a metric to gauge the model’s degree of surprise regarding the generated sequences. In this context, a lower perplexity value indicates a higher likelihood that the sequence has been encountered during training, thereby suggesting familiarity on the part of the model. Formally, perplexity is quantified as follows:
@@
where $log\quad f(x_{i}|x_{1},\ldots,x_{i-1},\Theta)$ indicates the log likelihood of the token $x_{i}$ given all previous tokens $x_{1},...,x_{i-1}$. Besides, a straightforward heuristic based on knowledge distillation can completely expunge the information that needs to be forgotten, while preserving an accuracy exceeding 90% on the retained data (chundawat2023zero).
The aforementioned studies collectively highlight that addressing the memorization problem of LLMs is crucial for safeguarding user privacy. Very recently, a latest study empirically demonstrates that the memorization problem also occurs in LLMs4Code, showcasing that hard-coded credentials in code can be easily leaked during the code completion process (huang2024your). Therefore, our study aims to understand how well the problem can be mitigated, which builds the foundation for mitigating the potential risks associated with the retention of sensitive information stored in LLMs4Code.
Figure 1: The workflow of our study.
@@ 2.3 Machine Unlearning
Relevant to the issue of privacy leaks in LLMs (huang2022large), the concept of unlearning involves recalibrating a model after training to erase certain contents from its captured knowledge (jagielski2020auditing; jang2022knowledge), thereby avoiding the costly process of retraining LLMs. Given the ability of large language models to recall specific details from their training sets, the targeted deletion of such privacy-sensitive data is of significant value.
Pioneer efforts by Chen & Yang (2023) (chen2023unlearn) and Eldan & Russinovich (2023) (eldan2023whos) have provided model designers with post-training modifications that can protect privacy at relatively low computational costs. Nevertheless, unlearning algorithms are still at an early stage of development, with different methods showing varying degrees of effectiveness and no definitive benchmark for evaluating their performance. To adequately assess the effectiveness of unlearning, a comprehensive evaluation framework is essential. This framework should include metrics that not only quantify the reduction of information related to the target data but also evaluate the functional behavior of the model after unlearning, ensuring that the overall performance of the model remains intact, while effectively omitting sensitive information. Our study aims at building such a benchmark for machine unlearning on LLMs4Code. Through rigorous evaluations on such a benchmark, we can move towards a more comprehensive understanding of the capabilities of machine unlearning and its significance in enhancing privacy-preserving practices within the application of LLMs4Code.
3 Study Design
Figure 1 demonstrates the workflow of this study. Suppose that LLMs4Code can initially work well on code generation (given the general query) but leak privacy information when given a specific code completion prompt (i.e., the privacy query). Our work aims to investigate whether applying various machine unlearning techniques to the models can (1) prevent sensitive information leakage in response to privacy-related queries, and (2) preserve code generation capabilities for general-purpose queries.
@@ 3.1 Research Questions
  • RQ3: Privacy Leakage Forms after Unlearning This question investigates how the sensitive information would still be leaked after unlearning. Especially, we focus on analyzing whether the leakage is intentional or unintentional.
  • RQ4: Fine-Grained Sensitive Information Leakage This study investigates how different unlearning techniques mitigate the sensitive information leaks across fine-grained data categories.
Figure 2: The prompt to transform the basic query to the code generation process.
@@ 3.2 Dataset
As for the forget set, we followed the previous study (maini2024tofu) and created a synthetic dataset consisting of 500 fictional character resumes. Each resume contains essential details such as name, address, education, phone number, and email address. The sensitive attributes in this study are categorized as follows:
@@
Figure 3: Prompt used to generate the basic information and the questions.
To evaluate the potential privacy risks posed by LLMs4Code, we designed 10 privacy-related questions for each fictional character. These questions are tailored to probe various aspects of personal data security and privacy concerns, including sensitive attributes like bank accounts and other distinct properties. In order to align with the main functionality of LLMs4Code (i.e., code generation), these questions are transformed into the format of code completions using GPT-4. The details are shown in Figure 2. In concrete terms, using GPT-4, the natural language query is converted into a Python function by extracting relevant information from the provided user ID card. This conversion involves generating code that represents the user’s details, such as name, occupation, and address, and outputs it in a function format. The transformation ensures that user data is structured in a way that facilitates automated code generation while maintaining context and privacy guidelines. Furthermore, these 5K questions are randomly split into train and test sets with a ratio of 9:1, where the train set is used to perform the unlearning, while the test set is used to evaluate the performance of the unlearning.
As for the retain set, we leveraged the well-known established datasets HumanEval (chen2021evaluating), MBXP (athiwaratkun2022multi), and LiveBenchwhite2024livebench in the code generation domain. Similar to the forget set, we randomly collected 5K samples from these two datasets and split them into train and test sets following a standard 9:1 ratio.
@@ 3.3 Unlearning Techniques
Forget Set Gradient Ascent (GA) (liu2022continual): The key idea of this approach is to calculate the gradients of the loss function with respect to the model’s parameters based on the responses to the privacy queries in the forget set. The gradients are then used to update the model’s parameters in a direction that encourages the model to ‘forget” the patterns associated with answering privacy questions directly. Formally,
@@
Forget Set Gradient Ascent and Retain Set Gradient Descent (GA+GD) (liu2022continual): The key idea of this approach is to simultaneously perform the forget set gradient ascent along with the retain set gradient descent, aiming at retaining the models’ capabilities to handle general queries. Formally,
@@
where $S_{R}$ denotes the retain set, $L(S_{R},w)$ is the loss function for the retain set $S_{R}$, and $L_{\text{diff}}$ represents the combined effect on the model’s loss.
Forget Set Gradient Ascent and Kullback-Leibler Divergence Minimization (GA+KL) (rafailov2024direct): This approach provides a dual-objective framework to simultaneously remove sensitive information and preserve general language modeling capabilities. While the gradient ascent on the forget set $S_{F}$ works to suppress the memorized private content, it may cause undesired side effects on the model’s generation ability. To mitigate this, we introduce a regularization term based on Kullback-Leibler (KL) divergence, which aligns the output distribution of the current unlearning model with that of a reference model fine-tuned only on the retain set $S_{R}$.
Concretely, we first obtain a clean model $M_{\text{fin}}$ via supervised fine-tuning (SFT) on the retain set. During the unlearning stage, we optimize the following loss:
@@
Here, $L(S_{F},w)$ denotes the cross-entropy loss over the forget set and is maximized (via gradient ascent) to actively disrupt memorized responses. The second term computes the token-level KL divergence between the probability distributions of the fine-tuned model $M_{\text{fin}}$ and the unlearning model $M_{\text{unl}}$ over each prefix $s_{<i}$ in the retain set. This encourages the unlearning model to retain similar generative behavior for non-sensitive inputs.
Compared to full loss minimization on the retain set (as in GA+GD), the KL term is lighter-weight and operates at the distribution level rather than over concrete ground-truth targets, making it a suitable regularization mechanism. This formulation enables the model to unlearn specific memorized information while preserving its broader knowledge and generation capabilities—especially in the auto-regressive decoding process of LLMs, where output distributions over partial sequences are crucial.
During training, both terms are jointly optimized: the gradient ascent on $S_{F}$ disrupts learned associations with privacy-related inputs, while the KL minimization on $S_{R}$ stabilizes the model’s behavior on general prompts, leading to a more balanced unlearning outcome.
Although the unlearning techniques we employ (e.g., gradient ascent and KL divergence) have been previously proposed and tested in natural language domains, their effectiveness in LLMs4Code remains largely unexplored. Unlike general LLMs, code models are often more sensitive to memorized tokens, such as function names, variable identifiers, or API patterns. This introduces new challenges in evaluating whether unlearning removes such memorization while maintaining functionality. Therefore, instead of relying on classification-based metrics, we design a dual-evaluation scheme focusing on privacy leakage and code generation accuracy to better reflect the characteristics of LLMs4Code.